The sum of the angles in each quadrilateral is 360 °. For example, in a rectangle 4 x 90 ° = 360 ° Gent writes: 5 x 90⁰ = 450⁰, so the sum of the angles in each pentagon must be 450 °.

Is he right? If not, give the correct answer.


1
Expert's answer
2021-05-27T06:53:46-0400

Unlike a rectangle, a Pentagon cannot have 5 right angles except they are all obtuse. A Pentagon can only have at most 3 right angles. Therefore, Gent's approach is wrong. To see this mathematically:

"\text{ Sum of interior angles= } (n-2) \times 180°"


Since Pentagon has 5 sides, n=5. Thus:

"=(5-2) \times 180°\\
=3 \times 180°\\
=540°"


Therefore, Gent is wrong.